The sum of the ages of Gerardo and Alexis is 104 years. 12 years ago, Gerardo's age was 3 times Alexis's age. How old is Gerardo now?

Found 5 solutions by ewatrrr, greenestamps, Solver92311, ikleyn, MathTherapy:
Answer by ewatrrr(24785)


Hi
Let x represent Alexis' age ..12 years ago
and 3x Gerardo's age ..12 years ago
12 years ago... 104 - 12 = 92
x + 3x = 92
4x = 92
x = 23 Alexis' age 12 years ago
23+12 = 35 Alexis' age NOW
35 + C(now) = 104
C(now)= 69 Gerardo's age NOW
CHECKING our answer***
35 + 69 = 104
Wish You the Best in your Studies.

Answer by greenestamps(13200)



I suspect the interpretation of the given information by the other tutor is not correct.

I think the information is that Gerardo's age 12 years ago was 3 times Alexis's age 12 years ago -- not 3 times Alexis's current age....

12 years ago, the sum of their ages was 104-2(12) = 80.

At that time, Gerardo was 3 times as old as Alexis. Any number of different methods can be used to determine that their ages 12 years ago were 60 and 20.

So Gerardo's age now is 60+12 = 72.


Answer by Solver92311(821)



Gerardo's age today:

Alexis' age today:

So:

Gerardo's age 12 years ago:

Alexis' age 12 years ago:

So: or more simply

Solve the 2X2 system for


John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by ikleyn(52798)

.
The sum of the ages of Gerardo and Alexis is 104 years. 12 years ago, Gerardo's age was 3 times Alexis's age.
How old is Gerardo now?
~~~~~~~~~~~~~~~~~


Let x be the Alexis' age.

Then the Gerardo age is  (104-x) years.



12 years ago

    Alexis age was (x-12) years,

    Gerardo's age was  (104-x-12) = (92-x),

        and the equation states

            92 - x = 3*(x-12)


Simplify it

    92 - x = 3x - 36

    92 + 36 = 3x + x

      128   = 4x

        x   = 128/4 = 32. 


ANSWER.  Gerardo's age is 104 - 32 = 72 years.

Solved.


Answer by MathTherapy(10552)


The sum of the ages of Gerardo and Alexis is 104 years. 12 years ago, Gerardo's age was 3 times Alexis's age. How old is Gerardo now?
Doesn't make sense, at all, to name a variable for one of the 2 people's ages 12 years ago
Having said that, let's make Gerardo's age, G
Then Alexis' is: 104 - G
We the get: G - 12 = 3(104 - G - 12)
G - 12 = 3(92 - G)
G - 12 = 276 - 3G
G + 3G = 276 + 12
4G = 288
Gerardo, or